# AI in Supply Chain Risk Assessment: A Systematic Literature Review and Bibliometric Analysis

### v2

@@ Abstract
Supply chain risk assessment (SCRA) has witnessed a profound evolution through the integration of artificial intelligence (AI) and machine learning (ML) techniques, revolutionizing predictive capabilities and risk mitigation strategies. The significance of this evolution stems from the critical role of robust risk management strategies in ensuring operational resilience and continuity within modern supply chains. Previous reviews have outlined established methodologies but have overlooked emerging AI/ML techniques, leaving a notable research gap in understanding their practical implications within SCRA. This paper conducts a systematic literature review combined with a comprehensive bibliometric analysis. We meticulously examined 1,717 papers and derived key insights from a select group of 48 articles published between 2014 and 2023. The review fills this research gap by addressing pivotal research questions and exploring existing AI/ML techniques, methodologies, findings, and future trajectories, thereby providing a more encompassing view of the evolving landscape of SCRA. Our study unveils the transformative impact of AI/ML models, such as Random Forest, XGBoost, and hybrids, in substantially enhancing precision within SCRA. It underscores adaptable post-COVID strategies, advocating for resilient contingency plans and aligning with evolving risk landscapes. Significantly, this review surpasses previous examinations by accentuating emerging AI/ML techniques and their practical implications within SCRA. Furthermore, it highlights the contributions through a comprehensive bibliometric analysis, revealing publication trends, influential authors, and highly cited articles. This analysis provides an encompassing view of the field’s growth, offering insights into its evolution. This comprehensive review provides a roadmap for practitioners and researchers, offering insights into fortifying supply chain risk management strategies through AI integration, ultimately contributing to a deeper understanding of evolving trends and applications in this dynamic field.
@@
Artificial Intelligence (AI), Supply chain risk assessment, Bibliometric analysis, Systematic literature review
@@
[inst1]organization=Department of Industrial Engineering and Management,addressline=Khulna University of Engineering and Technology (KUET), city=Khulna, postcode=9203, country=Bangladesh
@@
[inst2]organization=Department of Computer Science,addressline=American International University-Bangladesh (AIUB), city=Dhaka, postcode=1229, country=Bangladesh
@@ 1 Introduction
Supply chain management (SCM) has long grappled with various challenges stemming from a spectrum of sources. Past outbreaks of infectious diseases, geological upheavals like earthquakes, and other natural catastrophic events have put supply chains at risk, albeit on a limited scale (Govindan et al.,, 2020). These incidents have illustrated how interrelationships within supply chains have formed intricate risk contagion networks, making them susceptible to contagion effects (Agca et al.,, 2022). The cascade effect, characterized by risk spillover from one enterprise to another, exacerbates the challenges, culminating in a chain reaction of Supply Chain Resilience (SCR) disasters (Roukny et al.,, 2018).
Across the globe, companies confront formidable challenges across all stages of their supply chains. Suppliers failing to meet delivery obligations, unpredictable shifts in customer demands, and episodes of panic buying are just a few examples of the hurdles companies face (Ivanov,, 2020). In pursuing market leadership, enterprises must also contend with the complexity introduced by the management and launch of innovation projects, further complicating SCM (Kwak et al.,, 2018). Furthermore, digital transformation has ushered in a wave of technological advancements in supply chain operations (Kwak et al.,, 2018).
However, the most significant disruption in recent times emerged during the COVID-19 pandemic, impacting global supply chains profoundly. The disruptions encompassed not only the movement of people but also the flow of raw materials and finished goods, along with extensive disruptions in factory and supply chain operations (Sheng et al.,, 2021). These disruptions ushered in unprecedented challenges for supply chain professionals as they grappled with an entirely new reality (Araz et al.,, 2020; Craighead et al.,, 2020) . The pandemic represents a distinctive form of supply chain disruption, distinct from natural or man-made disasters (e.g., the Japanese earthquake and 9/11 attacks) and disruptions driven by evolving technologies and changing customer attitudes (Zhang et al.,, 2020).
@@
Supply chains are not the sole victims in this precarious environment, as the ripple effect extends to upstream and downstream enterprises. Practical risk management programs are imperative for enterprises to avert SCR and enhance SCR management (SCRM). Moreover, incorporating advanced AI technologies, such as ML, can provide predictive capabilities for SCRM. ML algorithms have showcased their ability to identify abnormal risk factors and derive predictive insights from historical data (Guo et al.,, 2021; Mohanty et al.,, 2021). By harnessing ML, enterprises can detect risk factors and anticipate market demands and potential risk scenarios (Punia et al.,, 2020; Wu et al., 2022a, ). ML’s proficiency in handling non-linear relationships further bolsters its superiority over traditional linear models. Additionally, its aptitude for processing unstructured data, a challenge where traditional models falter, positions ML as a formidable tool for addressing time, cost, and resource constraints within supply chains.
@@
However, despite its immense potential, supply chain researchers have historically exhibited limited familiarity with ML in comparison to other SCM aspects like mathematical programming and stochastic optimization (Liu et al.,, 2019; Shahed et al.,, 2021). Furthermore, the classification of ML algorithms remains unclear (Janiesch et al.,, 2021; Xu and Jackson,, 2019). Bridging this knowledge gap and exploring the value of ML for SCRM through interdisciplinary integration is a critical research need. Remarkably, there has been no prior effort to scrutinize the SCRM literature within an ML environment, in stark contrast to previous reviews primarily centered on risk definition, classification, and management strategies.
@@
Companies should augment their analytical capabilities by harnessing organizational knowledge to enhance SCR, thereby elevating their information capabilities (Wong et al.,, 2020). As underscored by previous studies, the role of AI-enabled technologies extends to promoting innovations for enhanced supply chain performance (SCP) (Baryannis et al., 2019b, ; Nayal et al.,, 2021). The adaptation capabilities and information processing prowess offered by AI techniques hold the potential to enhance SCP (Belhadi et al., 2021a, ). Notably, AI has found application across diverse sectors, improving flexibility and communication and reducing undue fluctuations for successful project execution (Lalmi et al.,, 2021). To navigate the impact of risks and disruptions, (Katsaliaki et al.,, 2022) advocate the integration of three critical facets: long-term partnerships, IT applications for business enhancement, and government policies that facilitate adaptability.
@@
The contemporary supply chain landscape demands the integration of AI, specifically ML, to invigorate SCRA and SCRM. While traditional methods possess merit, their limitations can be effectively addressed by AI, which offers predictive capabilities, nonlinear relationship analysis, and unstructured data processing. As supply chains continue to evolve, AI-based risk assessment is imperative for fostering resilience and sustaining the efficiency and effectiveness of supply chain operations. Existing reviews of SCRA using ML often fall short of offering a comprehensive view of evolving ML techniques, neglecting emerging trends and overlooking diverse applications and non-journal sources. They also tend to overlook the role of ML in the response phases of SCRM and do not sufficiently address data-related challenges.
@@
In light of the limitations associated with traditional SCRA methods, there is a growing interest in harnessing the potential of AI techniques to enhance risk assessment practices. This paper presents a systematic literature review (SLR) focused on the application of AI in SCRA. The main objective of this review is to analyze the existing literature critically, identify research gaps, and provide insights into the use of AI techniques, such as ML, deep learning(DL), and natural language processing(NLP), to improve the accuracy and effectiveness of SCRA. Our paper has the following contributions-
@@
    Our study significantly contributes to the literature by conducting a comprehensive review and employing bibliometric and cocitation analysis to assess the application of AI and ML techniques in SCRA, specifically focusing on the most recent papers.
@@
    This research addresses the evolving landscape of SCRA by considering the challenges posed by post-COVID uncertainties and emphasizing the increasing significance of AI in risk assessment.
@@
    This systematic review paper addresses the evolving landscape of SCRA by considering the challenges posed by post-COVID uncertainties and emphasizing the increasing significance of AI in risk assessment.
  4. 4.
    Our paper highlights pivotal trends and key findings, shedding light on the current landscape of AI/ML in SCRA.
  5. 5.
    We offer valuable insights into the next steps for researchers and practitioners, guiding the evolution of SCRA methods in an ever-changing world.
  6. 6.
    Our review aids academics and industry professionals seeking effective risk management strategies in today’s dynamic supply chain environment.
This review follows a structured paper skeleton, starting with an introduction to the topic and its significance in the field of SCM in the “Introduction” section. Following this, a comprehensive “Literature Review” unfolds, delving into the existing body of knowledge. The subsequent “Methodology” section outlines the study’s design, including article search strategies, selection criteria, and the synthesis of gathered data. Complementing these fundamental segments, the paper integrates a specific section for “Bibliometric Analysis” Utilizing bibliometric tools, this section showcases insights distilled from a comprehensive analysis of scholarly publications, focusing on AI in SCRA. Visualizations and key observations derived from this analysis are encapsulated here. Transitioning forward, the narrative navigates through diverse AI techniques deployed in SCRA, expounded within the “AI Techniques for Supply Chain Risk Assessment” section. Here, the discussion encapsulates the advantages, limitations, and real-world applications of these techniques. Moreover, it underlines the “Managerial Implications”, providing actionable insights and recommendations for industry practitioners based on the reviewed literature. This section delineates practical implications derived from the reviewed scholarly material aimed at aiding managerial decision-making within SCRA contexts. Continuing, the paper addresses the “Challenges and Limitations”encountered in this field, and forecasts potential “Future Research Directions” This forward-looking section paves the way for continued exploration and advancement within this domain. Ultimately, the review culminates in a comprehensive “Conclusion”, summarizing the primary findings drawn from the review process. Additionally, it furnishes recommendations tailored for both practitioners and researchers aimed at leveraging the full potential of AI in SCRA.
This meticulously structured approach endeavors to contribute to existing knowledge and bridge critical gaps within the literature. It is designed to furnish invaluable insights for academia and industry stakeholders alike.
@@ 2 Literature Review
Deiva Ganesh and Kalpana, (2022) conducted a comprehensive and descriptive literature study to identify AI and ML approaches in SCRM stages. This study examined SCRM research publications from three scientific databases from 2010 to 2021. They proposed a data analytics, simulation, and optimization framework that might produce a comprehensive Supply Chain risk identification, assessment, mitigation, and monitoring strategy. Hence, leveraging AI, Blockchain, and the Industrial Internet of Things (IIoT) to create smarter supply chains can transform how firms handle uncertainty. This analysis of (Deiva Ganesh and Kalpana,, 2022) does not evaluate blockchain, big data, and IIoT-related supply chain articles, which may limit information. Due to limited research, only recent English-language publications on SCRM and AI were included. Nimmy et al., (2022) did a thorough literature evaluation on operational risk assessment methods and whether AI explains them. The report suggested that risk managers use auditable supply chain operational risk management methodologies to understand why they should take a risk management action rather than just what to do. While they employed the SLR method, they could only examine some AI solutions for SCRM. They suggested hybridizing supply chain operational risk management and explainable AI (XAI) to obtain XAI-like characteristics in future research. Li et al., (2023) analyzed COVID-19 SCR research history, present, and future. In particular, supervised ML classifies 1717 SCR papers into 11 subject categories. Each cluster was then studied in the context of COVID-19, indicating three related skills (interconnectedness, transformability, and sharing) on which enterprises could work to develop a more resilient supply chain post-COVID. Their data only came from Scopus’ core collection, which might affect the results; thus, they suggested adding WoS and EBSCO to the evaluation Xu et al., (2020). Given the fast expansion of SCR research, they only picked English-language articles, which may have excluded valuable knowledge. They recommended network analysis to determine cluster linkages and SCR literary themes.
@@
Naz et al., (2022) examined the role of AI in building a robust and sustainable supply chain and offered optimal risk mitigation options. For review, 162 SCOPUS research publications were selected. Based on the nominated articles, Structural Topic Modeling was used to produce various AI-related theme topics in SCR. AI research trends in SCR were examined using R-package bibliometric analysis. They solely studied journal articles, not conference papers, field reports, corporate reports, book chapters, etc. Ni et al., (2020) analyzed articles from 1998/01/01 to 2018/12/31 in five major databases to highlight the newest research trends in the field. ML applications in SCM were still developing due to a lack of high-yielding authors and poor publication rates. 10 ML algorithms were extensively utilized in SCM, but their utilization was unequal among the SCM tasks most often reported. This paper has limitations in reviewing only five popular databases to limit articles for evaluation that might have filtered some related articles. Second, only widely used ML methods in SCM were counted in this review, and other ML techniques brought to SCM may be helpful later on. Low-frequency ML algorithms should be analyzed for further research in this area. Finally, their article contained 32 well-recognized ML methods; however, some newly generated ML algorithms may be used in SCM after 2015. Baryannis et al., 2019b examined supply chain risk definitions and uncertainty. Then, a mapping analysis categorizes available literature by AI approaches and SCRM tasks. Most of the examined works focus on building and assessing a mathematical model that accounts for various uncertainties and hazards but less on establishing and analyzing the applicability of the suggested models. They found that only 9 of 276 research (3%) use comprehensive techniques that cover all three SCRM stages (identification, assessment, and response). Yang et al., (2023) thoroughly examined the advancement of ML algorithms in SCRM by gathering 67 publications from 9 authoritative databases in the first half of 2021. They analyzed only English language journal articles from 9 relevant academic databases, excluding conference papers, textbooks, and unpublished articles and notes. This analysis relied on keyword searches, which may have missed some work. Schroeder and Lodemann, (2021) used a comprehensive and multi-vocal literature study to get a complete picture of our subject field. The SLR identified 533 papers in this area and examined 23. The comprehensive literature study found that just a few examples of ML in SCRM have been reported in depth in the scientific literature, and those examples focus on manufacturing, transport, and the complete supply chain. Shi et al., (2022) examined 76 works on credit risk utilizing statistical, ML, and DL methods over the last eight years. They offered a unique classification approach and performance rating for ML-driven credit risk algorithms utilizing public datasets. Data imbalance, dataset inconsistency, model transparency, and DL model underuse are discussed. Their review found that most DL models outperform standard ML and statistical algorithms in credit risk prediction, and ensemble techniques outperform single models.
Research Questions We formulated the following research questions to guide our SLR:

### v3

@@ Abstract
Supply chain risk assessment (SCRA) has witnessed a profound evolution through the integration of artificial intelligence (AI) and machine learning (ML) techniques, revolutionizing predictive capabilities and risk mitigation strategies. The significance of this evolution stems from the critical role of robust risk management strategies in ensuring operational resilience and continuity within modern supply chains. Previous reviews have outlined established methodologies but have overlooked emerging AI/ML techniques, leaving a notable research gap in understanding their practical implications within SCRA. This paper conducts a systematic literature review combined with a comprehensive bibliometric analysis. We meticulously examined 1,439 papers and derived key insights from 51 articles published between 2015 and 2024. The review fills this research gap by addressing pivotal research questions and exploring existing AI/ML techniques, methodologies, findings, and future trajectories, thereby providing a more encompassing view of the evolving landscape of SCRA. Our study unveils the transformative impact of AI/ML models, such as Random Forest, XGBoost, and hybrids, in substantially enhancing precision within SCRA. It underscores adaptable post-COVID strategies, advocating for resilient contingency plans and aligning with evolving risk landscapes. Significantly, this review surpasses previous examinations by accentuating emerging AI/ML techniques and their practical implications within SCRA. Furthermore, it highlights the contributions through a comprehensive bibliometric analysis, revealing publication trends, influential authors, and highly cited articles. This analysis provides an encompassing view of the field’s growth, offering insights into its evolution. This comprehensive review provides a roadmap for practitioners and researchers, offering insights into fortifying SCRM strategies through AI integration, ultimately contributing to a deeper understanding of evolving trends and applications in this dynamic field.
@@
Artificial Intelligence (AI) , Supply chain risk assessment , Bibliometric analysis , Content analysis , Systematic literature review
@@
[…]
[inst3]organization=Department of Computer Science,addressline=American International University-Bangladesh (AIUB), city=Dhaka, postcode=1229, country=Bangladesh
@@ 1 Introduction
Supply chain management (SCM) has long grappled with various challenges from various sources. Past outbreaks of infectious diseases, geological upheavals like earthquakes, and other natural catastrophes have put supply chains (SCs) at risk, albeit on a limited scale (Govindan et al.,, 2020). These incidents have illustrated how interrelationships within SCs have formed intricate risk contagion networks, making them susceptible to contagion effects (Agca et al.,, 2022). The cascade effect, characterized by risk spillover from one enterprise to another, exacerbates the challenges, culminating in a chain reaction of SC Resilience (SCR) disasters (Roukny et al.,, 2018).
Across the globe, companies confront formidable challenges across all stages of their SCs. Suppliers failing to meet delivery obligations, unpredictable shifts in customer demands, and episodes of panic buying are just a few examples of the hurdles companies face (Ivanov,, 2020). In pursuing market leadership, enterprises must also contend with the complexity introduced by the management and launch of innovation projects, further complicating SCM (Kwak et al.,, 2018). Furthermore, digital transformation has ushered in a wave of technological advancements in SC operations (Kwak et al.,, 2018).
However, the most significant disruption in recent times emerged during the COVID-19 pandemic, impacting global SCs profoundly. The disruptions encompassed not only the movement of people but also the flow of raw materials and finished goods, along with extensive disruptions in factory and SC operations (Sheng et al.,, 2021). These disruptions ushered in unprecedented challenges for SC professionals as they grappled with an entirely new reality (Araz et al.,, 2020; Craighead et al.,, 2020). The pandemic represents a distinctive form of SC disruption, distinct from natural or man-made disasters (e.g., the Japanese earthquake and 9/11 attacks) and disruptions driven by evolving technologies and changing customer attitudes (Zhang et al.,, 2020).
@@
SCs are not the sole victims in this precarious environment, as the ripple effect extends to upstream and downstream enterprises. Practical risk management programs are imperative for enterprises to avert SCR and enhance SC risk management (SCRM). Moreover, incorporating advanced AI technologies, such as ML, can provide predictive capabilities for SCRM. ML algorithms have showcased their ability to identify abnormal risk factors and derive predictive insights from historical data (Guo et al.,, 2021; Mohanty et al.,, 2021). By harnessing ML, enterprises can detect risk factors and anticipate market demands and potential risk scenarios (Punia et al.,, 2020; Wu et al., 2022a, ). ML’s proficiency in handling non-linear relationships further bolsters its superiority over traditional linear models. Additionally, its aptitude for processing unstructured data, a challenge where traditional models falter, positions ML as a formidable tool for addressing time, cost, and resource constraints within SCs.
@@
However, despite its immense potential, SC researchers have historically exhibited limited familiarity with ML compared to other SCM aspects like mathematical programming and stochastic optimization (Liu et al.,, 2019; Shahed et al.,, 2021). Furthermore, the classification of ML algorithms remains unclear (Janiesch et al.,, 2021; Xu and Jackson,, 2019). Bridging this knowledge gap and exploring the value of ML for SCRM through interdisciplinary integration is a critical research need. Remarkably, there has been no prior effort to scrutinize the SCRM literature within an ML environment, in stark contrast to previous reviews primarily centered on risk definition, classification, and management strategies.
@@
Companies should augment their analytical capabilities by harnessing organizational knowledge to enhance SCR, thereby elevating their information capabilities (Wong et al.,, 2020). As underscored by previous studies, the role of AI-enabled technologies extends to promoting innovations for enhanced SC performance (SCP) (Baryannis et al., 2019b, ; Nayal et al.,, 2021). The adaptation capabilities and information processing prowess offered by AI techniques hold the potential to enhance SCP (Belhadi et al., 2021a, ). Notably, AI has found application across diverse sectors, improving flexibility and communication and reducing undue fluctuations for successful project execution (Lalmi et al.,, 2021). To navigate the impact of risks and disruptions, (Katsaliaki et al.,, 2022) advocates integrating three critical facets: long-term partnerships, IT applications for business enhancement, and government policies that facilitate adaptability.
@@
The contemporary SC landscape demands the integration of AI, specifically ML, to invigorate SCRA and SCRM. While traditional methods possess merit, their limitations can be effectively addressed by AI, which offers predictive capabilities, nonlinear relationship analysis, and unstructured data processing. As SCs continue to evolve, AI-based risk assessment is imperative for fostering resilience and sustaining the efficiency and effectiveness of SC operations. Existing reviews of SCRA using ML often fail to offer a comprehensive view of evolving ML techniques, neglecting emerging trends and overlooking diverse applications and non-journal sources. They also tend to overlook the role of ML in the response phases of SCRM and do not sufficiently address data-related challenges.
@@
In light of the limitations associated with traditional SCRA methods, there is a growing interest in harnessing the potential of AI techniques to enhance risk assessment practices. This paper presents a systematic literature review (SLR) focused on the application of AI in SCRA. The existing literature lacks bibliometric analysis incorporating recent and adequate articles on the AI-SCRA topic. Another limitation is that most literature only contributed to bibliometric or content analysis. We tried to close this gap by including both analyses in our study, leveraging the systematic review methodology. The main objective of this review is to analyze the existing literature critically, identify research gaps, and provide insights into the use of AI techniques, such as ML, Deep Learning (DL), and natural language processing (NLP), to improve the accuracy and effectiveness of SCRA. Our paper has the following contributions-
@@
    Our study provides a comprehensive review of AI and ML applications in SCRA, supported by bibliometric and co-citation analyses to capture recent advancements and trends in the field.
@@
    We conduct content analysis to explore thematic trends, detailing the advantages, disadvantages, and specific application scenarios of various AI/ML models used in SCRA, offering insights into their suitability for different SC contexts.
@@
    Our proposed AI-based framework guides practitioners in assessing SC risks, from data collection to model selection and interpretation, facilitating efficient risk management in dynamic SCs.
The rest of the article is organized as follows: the “Literature Review” examines the current body of knowledge. The “Methodology” section details the study’s design, including search strategies and selection criteria. The “Bibliometric Analysis” provides insights from bibliometric tools, with visualizations and observations on AI in SCRA. A structured framework for SCRA is given in “Framework for SCRA using AI section. In the “Content Analysis of AI Techniques in SCRA,” the review discusses various AI techniques, highlighting their benefits, limitations, and real-world applications. Practical insights for industry are provided in the “Practical and Managerial Implications” section, offering actionable guidance for practitioners. The review also outlines key “Challenges and Limitations” in this area, and suggests “Future Research Directions” for ongoing exploration. The paper concludes with a summary of findings and recommendations in the “Conclusion.”
@@ 2 Literature Review
Deiva Ganesh and Kalpana, (2022) conducted a comprehensive and descriptive literature study to identify AI and ML approaches in SCRM stages. This study examined SCRM research publications from three scientific databases from 2010 to 2021. They proposed a data analytics, simulation, and optimization framework that might produce a comprehensive SC risk identification, assessment, mitigation, and monitoring strategy. Hence, leveraging AI, Blockchain, and the Industrial Internet of Things (IIoT) to create smarter SCs can transform how firms handle uncertainty. This analysis of Deiva Ganesh and Kalpana, (2022) does not evaluate blockchain, big data, and IIoT-related SC articles, which may limit information. Due to limited research, only recent English-language publications on SCRM and AI were included. Nimmy et al., (2022) did a thorough literature evaluation on operational risk assessment methods and whether AI explains them. The report suggested that risk managers use auditable SC operational risk management methodologies to understand why they should take a risk management action rather than just what to do. While they employed the SLR method, they could only examine some AI solutions for SCRM. They suggested hybridizing SC operational risk management and explainable AI (XAI) to obtain XAI-like characteristics in future research. Li et al., (2023) analyzed COVID-19 SCR research history, present, and future. In particular, supervised ML classifies 1717 SCR papers into 11 subject categories. Each cluster was then studied in the context of COVID-19, indicating three related skills (interconnectedness, transformability, and sharing) on which enterprises could work to develop a more resilient SC post-COVID. Their data only came from Scopus’ core collection, which might affect the results; thus, they suggested adding WoS and EBSCO to the evaluation (Xu et al.,, 2020). Given the fast expansion of SCR research, they only picked English-language articles, which may have excluded valuable knowledge. They recommended network analysis to determine cluster linkages and SCR literary themes.
@@
Naz et al., (2022) examined the role of AI in building a robust and sustainable SC and offered optimal risk mitigation options. For review, 162 SCOPUS research publications were selected. Based on the nominated articles, Structural Topic Modeling produced various AI-related theme topics in SCR. AI research trends in SCR were examined using R-package bibliometric analysis. They solely studied journal articles, not conference papers, field reports, corporate reports, book chapters, etc. Ni et al., (2020) analyzed articles from 1998/01/01 to 2018/12/31 in five major databases to highlight the newest research trends in the field. ML applications in SCM were still developing due to a lack of high-yielding authors and poor publication rates. 10 ML algorithms were extensively utilized in SCM, but their utilization was unequal among the SCM tasks most often reported. This paper has limitations in reviewing only five popular databases to limit articles for evaluation that might have filtered some related articles. Second, only widely used ML methods in SCM were counted in this review, and other ML techniques brought to SCM may be helpful later on. Low-frequency ML algorithms should be analyzed for further research in this area. Finally, their article contained 32 well-recognized ML methods; however, some newly generated ML algorithms may be used in SCM after 2015. Baryannis et al., 2019b examined SC risk definitions and uncertainty. Then, a mapping analysis categorizes available literature by AI approaches and SCRM tasks. Most of the examined works focus on building and assessing a mathematical model that accounts for various uncertainties and hazards but less on establishing and analyzing the applicability of the suggested models. They found that only 9 of 276 research (3%) use comprehensive techniques that cover all three SCRM stages (identification, assessment, and response). Yang et al., (2023) thoroughly examined the advancement of ML algorithms in SCRM by gathering 67 publications from 9 authoritative databases in the first half of 2021. They analyzed only English language journal articles from 9 relevant academic databases, excluding conference papers, textbooks, and unpublished articles and notes. This analysis relied on keyword searches, which may have missed some work. Schroeder and Lodemann, (2021) used a comprehensive and multi-vocal literature study to get a complete picture of our subject field. The SLR identified 533 papers in this area and examined 23. The comprehensive literature study found that just a few examples of ML in SCRM have been reported in depth in the scientific literature, and those examples focus on manufacturing, transport, and the complete SC. Shi et al., (2022) examined 76 works on credit risk utilizing statistical, ML, and DL methods over the last eight years. They offered a unique classification approach and performance rating for ML-driven credit risk algorithms utilizing public datasets. Data imbalance, dataset inconsistency, model transparency, and DL model underuse are discussed. Their review found that most DL models outperform standard ML and statistical algorithms in credit risk prediction, and ensemble techniques outperform single models.
Previous studies in SCRM have shown limitations in evaluating certain technologies, such as blockchain, big data, and the IIoT, potentially limiting the scope of analysis. Our review addresses this gap by incorporating an analysis of these technologies within the context of SCRM, providing a more comprehensive assessment. There is a recognized need in the literature to integrate SC operational risk management with XAI for future research. Our review contributes to closing this gap by exploring the integration of these approaches and discussing their implications for SCRM. Some studies have highlighted the importance of utilizing multiple databases to evaluate SCRM literature comprehensively. Our review extends this by considering insights from the most reliable and prominent two databases, enhancing the breadth of our analysis. Previous research has found that only a small percentage of studies cover all three stages of SCRM comprehensively. Our review contributes to closing this gap by providing a holistic assessment of SCRM stages, including identification, assessment, and response strategies, leveraging insights from diverse studies.
3 Methodology
3.1 Study Design
The research utilized a systematic approach to conduct a comprehensive literature review and a bibliometric analysis. This methodology aimed to scrutinize the intersection of AI/ML techniques and SCRA, providing insights into research trends, methodologies, and emerging themes. We formulated the following research questions to guide our SLR: